Help please....Some brands of grapefruit juice are only 10% juice. Other brands are 40% juice. How much of each kind of juice should be mixed together to form 5 gallons of punch which is 25% juice?
I've tried this one several times and am confused. Thank You

Answer by mukhopadhyay(490)

Assume x gallons of 10% juice is to be mixed with (5-x) gallons of 40% juice to produce 5 gallons of 25% juice.
=> 10% of x + 40% of (5-x) = 25% of 5
=> .1x + .4(5-x) = .25(5)
=> .1x + 2 - .4x = 1.25
=> .3x = .75
=> x = .75/.3 = 2.5
Answer: 2.5 gallons of 10% juice is to be mixed with 2.5 gallons of 40% juice to produce 5 gallons of 25% juice.